The distance x covered by a particle varies with time t as x2=2t2+6t+1. Its acceleration varies with x as:
1.x
2. x2
3. x-1
4. x-3

Grade:11

1 Answers

Khimraj
5 years ago
x2=2t2+6t+1
x = (2t2+6t+1)3/2
differentiating wrt t
v = (2t+3)/(2t2+6t+1)1/2
angain differentiating wrt t
a = -7/(2t2+6t+1)3/2 = -7/x3
so a \propto x-3
Hope it clears...........................